manufacturer does not know the mean and SD of the diameters of ball bearings he is 

producing. However, a sieve system rejects all bearings larger than 2.4cm and those under 

1.8cm in diameter. Out of 1000 ball bearings 8% are rejected as too small and 5.5% as too 

big. What is the mean and standard deviation of the ball bearings produced assuming that 

the diameters of ball bearings are normally distributed?


1
Expert's answer
2022-05-16T15:37:57-0400
"P(X<x_1)=P(Z<\dfrac{x_1-\mu}{\sigma/\sqrt{n}})"

"=P(Z<\dfrac{1.8-\mu}{\sigma/\sqrt{1000}})=0.08"

"\dfrac{1.8-\mu}{\sigma/\sqrt{1000}}=-1.40507"


"P(X>x_2)=1-P(X\le x_2)"

"=1-P(Z\le\dfrac{x_2-\mu}{\sigma/\sqrt{n}})"

"=1-P(Z\le\dfrac{2.4-\mu}{\sigma/\sqrt{1000}})=0.055"

"\dfrac{2.4-\mu}{\sigma/\sqrt{1000}}=1.59819"

"\dfrac{2.4-\mu}{1.8-\mu}=\dfrac{1.59819}{-1.40507}"

"-3.372168+1.40507\mu=2.876742-1.59819\mu"


"\mu=2.0807"

"\dfrac{2.4-2.0807}{\sigma/\sqrt{1000}}=1.59819"

"\sigma=6.318"